### v2

# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

Abstract
The escalating scale and cost of Large Language Models (LLMs) training necessitate accurate pre-training prediction of downstream task performance for efficient resource allocation. This is challenged by: 1) the emergence phenomenon, where metrics become meaningful only after extensive training, hindering prediction by smaller models; and 2) uneven task difficulty and inconsistent performance scaling patterns, leading to high metric variability. Current prediction methods lack accuracy and reliability. We propose a Clustering-On-Difficulty (COD) framework for downstream performance prediction. The COD framework clusters tasks by their difficulty scaling features, thereby establishing a more stable and predictable support subset through the exclusion of tasks exhibiting non-emergent behavior or irregular scaling. We adopt a performance scaling law to predict cluster-wise performance with theoretical support. Predictable subset performance acts as an intermediate predictor for the full evaluation set. We further derive a mapping function to accurately extrapolate the performance of the subset to the full set. Applied to an LLM with 70B parameters, COD achieved a 1.36% average prediction error across eight key LLM benchmarks, offering actionable insights for resource allocation and training monitoring of LLMs pretraining.
[…]
Chenyin Xu at , Ke Shen at
1 Introduction
Large Language Models (LLMs) have emerged as transformative technologies in natural language understanding, generation, and reasoning [1, 15, 5]. Their impressive success heavily relies on scaling model parameters and pre-training data, with training loss empirically following a power-law relationship with compute [19, 22]. However, this reduction in training loss primarily reflects an in-domain compression effect and does not necessarily indicate improved out-of-domain generalization or downstream performance–the factor of primary concern in practice. Specifically, performance scaling of downstream tasks aims to predict the accuracy of the target LLM on downstream tasks using metrics from smaller models. Our objective is to develop a prediction method that works reliably on a diverse range of downstream tasks, optimizing the worst-case prediction error.
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [6, 13, 8, 38, 28], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [43]. Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 20], yet the uneven difficulty distribution between different evaluation samples undermines its accuracy.
A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give a detailed analysis in Sec. 3.
To address these challenges, we propose a new performance scaling law, derived from the existing loss scaling law [22], specifically applicable to evaluation subsets that exhibit consistent performance scaling patterns. Building on the performance scaling law, we develop a Clustering-On-Difficulty (COD) multi-stage framework for predicting downstream performance. Specifically, we first filter out clusters that lack scaling properties using an improved MeanShift clustering algorithm. Next, we fit the performance-compute relationships in the remaining clusters under our performance scaling law, extrapolate the performance of large models within these clusters, and finally map the aggregated predictions to the complete task set.
We validate our COD approach on eight popular evaluation sets, including MATH [16], BBH [33], and MMLU pro [36] datasets. COD achieves an average prediction error of 1.36% on an LLM with 70B parameters. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
Our contributions can be summarized as follows.
  • We propose the COD framework to address high variance and emergent phenomena in LLM performance scaling by effectively modeling the difficulty distribution within the evaluation sets.
  • We introduce a performance scaling law for cluster-wise performance prediction, with theoretical support and experimental validation.
  • Extensive experiments conducted in eight different evaluation sets demonstrate that COD provides reliable predictions with an average prediction error of 1. 36% on an LLM with 70B parameters.
2 Related Work
2.1 Loss Scaling Laws
Loss scaling laws provide a systematic framework for understanding the relationship between computational resources, data, model size, and the LLM performance. Early work by Kaplan et al. [22] demonstrates that the pre-training loss of LLMs follows a power-law relationship with the compute (the number of floating-point operations) used in training. Subsequent studies extend these findings to other domains, such as computer vision [42], graph learning [26], and vision-language models [2, 17]. Recent research has also explored scaling laws in specific contexts, such as fine-tuning [18, 35], vocabulary size optimization [34], retrieval-augmented models [31], and hyperparameter tuning [24, 41]. These studies highlight the broad applicability of scaling laws and their potential to guide the efficient allocation of computational resources.
2.2 Downstream Task Performance Scaling
Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that manifest only after exceeding task-specific thresholds [37, 29]. Recent works, such as using loss as a proxy [6] or increasing metric resolution [20], have demonstrated potential but encounter challenges in aligning surrogate metrics with original task objectives. Here, we briefly review the two main types of methods for predicting downstream performance:
Loss-intermediate prediction. These methods predict the final training loss (or in-domain validation loss) of LLMs with loss scaling laws first, and then predict downstream performance through loss-performance relationships [6, 13, 8]. While these methods leverage established scaling laws for loss predictions, they encounter a fundamental limitation: the inconsistent mapping between loss and performance metrics. In addition, Xiao et al. [38] employ the evaluation set answer loss as an intermediate variable for estimation. Although answer loss correlates with the final performance metrics, its predictability remains low as predicting answer loss shares the challenges with predicting performance, including emergence phenomenon and high variance in task difficulty.
End-to-end performance-compute prediction. These methods [20, 28, 1] directly model the relationship between performance and compute (or the number of model parameters). Additionally, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set. Hu et al. [20] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
3 Pilot Study
Figure 1: Performance-loss relationship across different model sizes (left) and learning rate schedules (middle). Performance-compute relationship for different clusters of the BBH samples(right)
In this section, we present the pilot experiments to illustrate the shortcomings of existing approaches.
Training loss may mismatch downstream task performance. Predicting downstream performance from training loss assumes LLMs achieve identical downstream results at the same loss value, which does not hold universally. In practice, training loss primarily serves as an indicator of in-domain fitting, whereas downstream tasks typically represent out-of-domain evaluations. Moreover, training configurations, such as model size and learning rate, can significantly affect not only the final loss but also the model’s generalization capabilities.
Fig. 1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [30]. At the same training loss level, smaller models can outperform larger ones in terms of test accuracy. Because smaller models initially exhibit weaker in-domain fitting capacity, they typically require more training steps to reach the same loss value, which can lead to better in-domain generalization once they do. Fig. 1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [7]. At the same loss level, the performance under the cosine schedule is always worse than that under the constant schedule, indicating that a lower learning rate may prioritize memorization over generalization, thereby diminishing downstream performance.
Diverse scaling patterns within the evaluation set. Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples. Fig. 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [33], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
Taken together, these observations highlight the importance of modeling the heterogeneous scaling properties within an evaluation set and identifying a robust intermediate metric to serve as a reliable indicator of the downstream performance of LLMs.
Figure 2: The pipeline of Cluster-On-Difficulty downstream task performance scaling, including 4 stages: a. Represent task difficulty feature with task-wise passrate vector. Cluster on the difficulty feature and filter outliers. b. Fit cluster-wise performance-compute curve. Classify clusters into extrapolatable clusters, non-extrapolatable clusters, and non-emergent clusters. c. Predict accuracy on extrapolatable clusters. d. Map subset accuracy prediction to full evaluation set performance.
4 Method
In this section, we introduce the COD method in four parts, as illustrated in Fig. 2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Sec. 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec. 4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Sec. 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec. 4.4).
4.1 Clustering on Difficulty
Tasks within evaluation sets, despite common themes, show substantial difficulty differences, leading to diverse scaling patterns that hinder universal fitting functions. We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster. To quantify task difficulty, we leverage the passrate(mean accuracy across multiple trials) from small models of progressively increasing scales. This process generates a feature vector for each task, which ideally exhibits monotonic growth, thereby reflecting the incremental enhancement of model capabilities.
Improved clustering methods. We hope to adopt clustering algorithms with the following features: 1. Minimizing intra-class variance to ensure similar extrapolation properties within each cluster, 2. Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
While K-Means [27] requires pre-specifying cluster numbers and DBSCAN [11] may yield large intra-cluster variance, MeanShift [12] allows constraining this variance via a radius parameter, better suiting our needs.
Figure 3: t-SNE visualization of different clustering methods: DBSCAN(left), MeanShift(Middle), Improved-MeanShift(Right). Each point represents an evaluation sample.
To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of evaluation tasks on BBH [33]. Each point represents an evaluation sample, and its color denotes the cluster type. As shown in Fig. 3, our improved MeanShift effectively splits dense areas. We explain the details of clustering algorithms in Sec. A.1, and smoothing techniques in Sec. A.2.
4.2 Fitting
After clustering, we compute metrics for small models within each cluster. These small models are trained with the same ratio of training tokens to compute per token. We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers. We derive the following fitting formula for the downstream task scaling law based on the training loss scaling law.
Theorem 1 (Scaling Law for Downstream Task Performance).
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
  1. 1.
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
  2. 2.
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
  3. 3.
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (1)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
Proof Sketch.
Assumption 3 (unique answers, no pre-answer reasoning) implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and individual task passrate, as correct task completion probability becomes the product of answer token probabilities. A key distinction, often overlooked, is how averaging occurs: the passrate metric averages $\exp(-\mathrm{loss}_{\mathrm{ans}})$ (an arithmetic mean), whereas applying the loss scaling law effectively averages loss before exponentiation (yielding a geometric mean). We show that the difference between these means can be approximated by $\frac{\sigma^{2}}{\mu}$ (where $\sigma^{2}$ and $\mu$ are the variance and mean of task passrates, respectively). Consequently, a performance scaling law derived from the loss scaling law is most valid for evaluation sets with limited difficulty variance. Our COD method achieves this by clustering tasks, thereby constraining intra-cluster difficulty variance and enabling better alignment with the performance scaling law. Finally, for tasks with a finite answer set, a random guessing score $g$ accounts for expected performance on unmastered tasks. Detailed proofs are provided in Appendix B. ∎
theorem 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$y(C)=g+(1-g)*e^{-aC^{-b}-c},$ (2)
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster. Parameters $a$, $b$, $c$, and $g$ are to be fitted. Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix G.
4.3 Extrapolation
To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction.
A cluster is deemed extrapolatable if it meets two criteria: (1) its expected accuracy increases monotonically with model size, and (2) its performance converges to at least a predefined threshold $P$ (where $P\leq 1$ accounts for practical limits like ambiguous questions or finite training coverage).
We filter out non-extrapolatable clusters using two rules based on the parameters from Eq. 2:
  1. 1.
    Negligible accuracy growth, indicated by minimal $a$ or $b$ values.
  2. 2.
    Poor extrapolation reliability, indicated by an excessive $c$ value.
In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $10\leq c<$ 1.
The clusters satisfying these conditions form the predictable subset. The final performance prediction for a target model on this subset is the weighted average of the extrapolated predictions from these individual clusters, with weights proportional to cluster sizes.
4.4 Mapping from Predictable Subset to Target Evaluation Set
We map predictions from the predictable subset $T^{\prime}$ to the complete evaluation set $T$ using a function $f$. This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering. The mapping function $f:T^{\prime}\rightarrow T$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$. Empirical validation indicates a quartic function optimally captures this relationship:
$f(x)=\alpha_{1}x^{4}+\alpha_{2}x^{3}+\alpha_{3}x^{2}+(1-\alpha_{1}-\alpha_{2}-\alpha_{3})x$ (3)
To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors. This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [40]) as anchors for many tasks (see Sec. 5.3.3 for experiments). For data-sensitive tasks, however, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy. The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq. 2 with the mapping $f$.
5 Experiments
5.1 Experimental Setups
In our experimental setup, we train several smaller versions of the target model architecture for prediction. These models vary in size, but share similar training procedures, with the training data scaled proportionally to their sizes. We show the training configurations of the model in Appendix C.
For evaluation, we adopt the following widely used benchmarks, including GSM8K [7], MATH [16], BBH [33], TriviaQA [21], MBPP [3], AGIEval [44], DROP [9], MMLU-pro [36]. All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [10]. We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks. Using a series of smaller models with identical data distribution and architecture but varied configurations, we estimate the downstream task performance of a pretrained LLM with 70B parameters.
5.2 Prediction Experiments
We compare COD against three representative prediction methods:
  1. 1.
    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws.
  2. 2.
    Passrate-compute prediction [1, 20]: Estimates large model passrates from smaller model passrates. We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
  3. 3.
    Loss-intermediate performance prediction [6]: First predicts the target LLM’s final training loss, then estimates downstream task metrics based on the relationship between smaller models’ evaluation metrics and their losses.
Table 1: Absolute prediction error (in percentage points) on evaluation sets. Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red). $\downarrow$ indicates lower is better.
Method Overall Metrics Individual Task Sets
[…]
End-to-end 3.10 6.00 4.00 3.86 0.64 0.68 1.75 6.00 4.11 3.72
Passrate 5.02 8.80 6.71 8.80 3.51 4.00 7.34 6.78 0.26 2.74
Loss-intermediate 5.29 9.39 9.39 6.95 2.33 5.81 5.52 1.41 5.37 5.55
COD (w/o mapping) 2.24 5.26 4.70 0.50 2.91 1.98 0.89 5.26 1.08 0.57
COD (Complete) 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
We also evaluate two variants of our COD approach to validate the benefits of its components:
  1. 1.
    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within each cluster, and aggregates metrics, bypassing the subset-to-full mapping stage.
  2. 2.
    COD (Complete): Our full proposed multi-stage approach, including clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
Comparative results are shown in Table 1. Prediction accuracy is measured by the absolute error between predicted and actual performance. We report mean and max prediction errors across all evaluation sets, as well as errors for individual sets. Our complete COD approach significantly outperforms existing methods in both mean (1.68%) and maximum (2.38%) prediction errors, offering reliable guidance for large model training. While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
Figure 4: Performance-compute relationship for different prediction methods on BBH, MATH, and MMLU-pro evaluation sets.
Fig. 4 visualizes the performance-compute relationships. On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly. COD reveals a more complex, better-fitted multi-phase trajectory. On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error. COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets (e.g., MATH) and capture diminishing returns in saturated sets.
The COD method is also capable of predicting LLMs’ performance after continual training phase on high-quality data. We conduct experiments and analysis on continual-trained LLMs and report prediction results in Appendix D.
5.3 Ablation Study
5.3.1 Comparison of Clustering Methods
We assess the impact of different clustering algorithms on prediction accuracy. The goal is to achieve tight intra-cluster difficulty similarity (low average distance to center) while maintaining cluster stability (min. 10 samples/cluster). We compare our Improved-MeanShift with DBScan, MeanShift, and K-Means. For K-Means, we adjust it to approximate our goals: (1) search for $k$ yielding min. cluster sizes $\approx 10$; (2) treat samples outside a radius (e.g., 2x average intra-cluster distance) from any cluster center as outliers, ensuring clusters don’t drop below 10 samples. We term this "Improved-KMeans" for this comparison. Clustering quality is measured by Intra-cluster Average Distance (IAD) and Outlier Rate (OR) (Table 2). Prediction benefits are measured by Extrapolation Error (EE) on the predictable subset and Final prediction Error (FE) (Table 3).
[…]
Method MMLU-pro GSM8k MATH BBH
[…]
K-Means 0.3236 - 0.2238 - 0.2238 - 0.6284 -
DBScan 0.4242 0.56 0.5131 0.53 0.4775 0.68 0.7113 18.92
MeanShift 0.2859 0.39 0.2852 0.61 0.2110 1.44 0.2679 20.72
Improved-KMeans 0.1609 2.85 0.1321 2.73 0.0902 2.22 0.1953 37.23
Improved-MeanShift 0.2225 4.40 0.1854 4.93 0.1463 2.66 0.2143 33.58
[…]
Method Mean Max MMLU-pro GSM8k MATH BBH
[…]
K-Means 3.62 3.76 8.16 8.99 3.69 3.69 0.01 0.00 2.62 2.34 8.16 8.99
DBScan 3.93 4.08 4.38 4.36 3.72 3.69 4.08 4.12 4.38 4.16 3.53 4.36
MeanShift 2.12 1.68 3.15 3.08 3.15 3.08 0.67 0.74 2.55 2.26 2.12 0.65
Improved-KMeans 1.33 1.84 3.92 4.08 0.56 0.61 3.92 4.08 0.81 0.51 0.02 2.17
Improved-MeanShift 1.23 1.66 2.20 2.23 1.27 1.35 2.20 2.23 1.14 1.28 0.29 1.77
Table 2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints. Table 3 confirms these methods lead to lower EE and FE. Although Improved-KMeans has the best IAD, its GSM8k prediction is poor. This is likely because K-Means requires pre-specifying $k$, and our search for $k$ can be unstable, leading to large errors on some sets. In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error. Additional results in Sec. E.1 support these conclusions.
5.3.2 Different Extrapolation Formulas
Table 4: Ablation study of extrapolation formulas. EE, TR, FE shown for BBH, MATH, MMLU-pro.
Method BBH MATH MMLU-pro
[…]
Direct Power Law ($f_{3}$) 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
w/o Random Guess ($f_{1}$) 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
w/o Constant c ($f_{2}$) 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
Ours ($f$) 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: 1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); 2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); 3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [20]). Table 4 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE). Our proposed formula $f$ consistently achieves the lowest EE and FE. $f_{1}$ struggles with finite-answer tasks where small models have non-zero scores. $f_{2}$ inaccurately assumes perfect scores are attainable, ignoring data limitations and task ambiguities. The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation. TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping (difficulty distributions visualized in Appendix F).
5.3.3 Different Anchor Points in Interpolation Mapping
We find that the mapping relationship from predictable subset metrics to full evaluation set metrics is similar across models with different training data and architectures. This allows leveraging pre-trained models as "anchors" to refine the mapping and improve final estimation accuracy. We use Qwen2-72B [40] (OOD anchor) and an in-house MoE model [23] with the same training data but different architecture (ID anchor). We first derive an interpolation curve using only small model metrics and fixed points $(0,0)$, $(1,1)$, then assess anchor compatibility. This shared mapping implies estimable subset metrics are highly correlated with full-set metrics and less prone to interference from other model parameters than loss-intermediate predictions.
We test three configurations:
  • COD (w/o anchor): Full COD pipeline, but no anchor points in the mapping phase;
  • COD (w. OOD anchor): COD with Qwen2-72B as an out-of-distribution anchor;
  • COD (w. ID anchor): COD with our in-house MoE model as an in-domain anchor.
Table 5: Influence of anchor points in the mapping stage on prediction error (pp). Lower is better.
Method Overall Individual Task Sets
[…]
w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
w. OOD anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
Table 5 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy. This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models. Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models. Additional ablations on the interpolation method (Sec. E.2) confirms quartic functions are suitable.
6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions. We discuss the limitations of the COD method in Appendix G. Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
Boarder Impact
The LLMs performance scaling prediction method proposed in this work can be used to guide decision-making for training LLMs and to monitor whether the LLM training process aligns with expectations. Accurate prediction allows for better planning of compute and human resources, thereby playing a valuable role in scientifically advancing the training of larger LLMs. However, if the prediction results are inaccurate, it may lead to outcomes that do not meet expectations for the invested resources, potentially resulting in resource waste or economic losses.
Appendix A Improvements of Clustering Algorithm
A.1 Improved MeanShift Algorithm
We iteratively apply the MeanShift algorithm with a predefined cluster radius $R$ and a minimum cluster size $K$. In each iteration, for the clustered samples, we examine whether the distance between each sample and its cluster center exceeds R, and relabel those samples that exceed this threshold as unclustered. For clusters containing fewer than K samples, we mark all samples in these clusters as unclustered. At the end of each iteration, we incorporate both the outliers from MeanShift and our marked unclustered samples into the next round of clustering, continuing this process until no further changes occur in sample labels. We present the pseudocode in algorithm 1.
Algorithm 1 Iterative MeanShift Clustering Algorithm
1:Initialize all labels in the evaluation set to $-1$
2:repeat
3:     Perform MeanShift clustering with radius $R$ on all samples labeled $-1$
4:     Assign new labels to clustered samples
5:     for each newly labeled sample $i$ do
6:         Calculate distance $dist_{i}$ to its cluster center
7:         if $\mathrm{dist}_{i}>R$ then
8:              Reset label to $-1$
9:         end if
10:     end for
11:     for each cluster do
12:         if number of samples in cluster $<K$ then
13:              Reset all samples in this cluster to $-1$
14:         end if
15:     end for
16:     Renumber all non-$\{-1\}$ newly labeled samples to avoid overlap with old labels
17:until no label changes
Filtering zero-performance samples.
In the evaluation set, there may exist a few extremely difficult problems that require sufficient model parameters to emerge. All small models may fail to solve these problems even after 100 evaluation attempts, resulting in difficulty feature vectors of all zeros. We refer to these as zero-performance samples. Their presence leads to two issues:
  1. 1.
    Zero performance on small models does not necessarily indicate zero accuracy on large models. For these samples, we cannot estimate when emergence will occur or predict large model metrics.
  2. 2.
    During clustering, they may be confused with other low-performing but non-zero samples. Including them in the same cluster would lower the expected accuracy of that cluster, leading to inaccurate fitting and extrapolation later.
Therefore, we pre-filter these zero-performance samples before clustering, treating them as outliers that do not participate in the clustering process. This approach obviates the necessity of considering their metrics under large models during subsequent extrapolation, and prevents disruption to the clustering of the remaining samples.
A.2 Smoothing Techniques
Metric fluctuations of individual samples in downstream tasks are not solely due to limited sampling. Another potential factor is noise from uneven data distribution in recent training batches. Therefore, in addition to performing 100 evaluations to mitigate sampling variance, we evaluated 100 times on each of the adjacent checkpoints before and after the selected model. We then averaged these accuracy expectation values across three checkpoints, further reducing sampling variance while offsetting noise from uneven training data distribution. This approach also reduces the number of zero-performance samples, further improving clustering and prediction effectiveness.
Appendix B Proof of Theorem 1
We use lemma 1 to derive the scaling law for downstream task performance (theorem A1).
Lemma 1 (Arithmetic-geometric mean difference).
For any sequence of positive real numbers $\{x_{i}\}_{i=1}^{n}$, let:
  • $\mu_{a}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the arithmetic mean;
  • $\mu_{g}=\prod_{i=1}^{n}x_{i}^{1/n}$ be the geometric mean;
  • $\sigma^{2}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu)^{2}$ be the variance.
Then the difference between the arithmetic mean and geometric mean can be estimated as:
$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a}).$ (4)
Proof.
Taking the logarithm of the geometric mean $\mu_{g}$:
$\log(\mu_{g})=\frac{1}{n}\sum_{i=1}^{n}\log x_{i}$ (5)
Using Taylor expansion of $\log x$ around $\mu$:
$\log x=\log\mu+\frac{x-\mu}{\mu}-\frac{(x-\mu)^{2}}{2\mu^{2}}+o\left((x-\mu)^{2}\right).$ (6)
We can simplify:
$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
$=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o(\mu_{a})$
$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a}).$
Therefore:
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a}).$ (7)
When $\frac{\sigma^{2}}{2\mu^{2}}$ is small, this can be approximated as:
$\Delta\approx\frac{\sigma^{2}}{2\mu_{a}}$ (8)
Theorem A1 (Scaling Law for Downstream Task Performance).
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
  1. 1.
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
  2. 2.
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
  3. 3.
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (9)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
Proof.
We first use Assumption 3 to establish the relationship between model passrate and loss on a task:
$-\log(p_{\mathrm{ans}})$ $=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}.$ (10)
Then take the exponential of both sides, and then take the expectation with respect to different tasks in the evaluation set $p=(q,\mathrm{ans})\in P$. We note that both $p_{\mathrm{ans}}$ and $\mathrm{loss}_{\mathrm{ans}}$ are functions of $C$.
$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (11)
We can adopt lemma 1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
$\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\mathrm{loss}_{\mathrm{ans}}(C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (12)
$=$ $\exp{(-aC^{-b}-c)}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (13)
where $n=\#P$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
Finally, we use assumption 2 to align the answer passrate and the accuracy metric. We can adopt the law of total expectation:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=$ $\mathbb{P}_{p}(\mathrm{correct})\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]+\mathbb{P}_{p}(\mathrm{incorrect})\mathbb{E}[\mathrm{Acc|incorrect}]$ (14)
Note that $\mathbb{P}_{p}(\mathrm{correct})=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$, $\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]=1$ and $\mathbb{P}_{p}(\mathrm{incorrect})=1-\mathbb{P}_{p}(\mathrm{correct})$. We also define $g$ as the random guess accuracy performance, thus we have $\mathbb{E}[\mathrm{Acc|incorrect}]=g$. Take these results into Eq. 14, and we have:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]+(1-\mathbb{E}_{p}[p_{\mathrm{ans}}(C)])*g$ (15)
$=g+(1-g)\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ (16)
$=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (17)
theorem A1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}.$ (18)
Appendix C Experiment Settings
Model training. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters. All models are trained on a constant learning rate scheduler with a few-step warmup stage and data with the same distribution. We list model configurations in Table A1.
Table A1: Model architecture specifications across different sizes.
122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
Compute Per Token (B) 1.535 2.684 4.275 6.378 9.060 16.436 54.761 91.609 475.131
Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
Continue-Trained Tokens (B) 3 5 8 12 18 33 114 191 1,000
Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
FFN Dimension 3,584 4,480 5,376 6,272 7,168 8,960 14,336 16,128 28,672
Heads 8 10 12 14 16 20 32 36 64
KV Heads 8 10 12 14 16 20 8 12 8
Evaluation settings. We conducted performance scaling estimation experiments across 8 major LLM evaluation sets. These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities. All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics. Detailed information is provided in Table A2.
Table A2: Information of evaluation datasets.
Dataset Domain #Questions #Shots
GSM8K Math 1,319 8
MATH Math 5,000 4
BBH Reasoning 6,511 3
TriviaQA Knowledge 17,944 5
MBPP Coding 500 3
AGIEval Comprehensive 8,063 5
DROP Reading 9,536 3
MMLU-pro Comprehensive 12,032 5
Appendix D Performance Prediction for Continue-Pretrained LLMs
Leading industry pre-trained LLMs (e.g., Deepseek-v3 [25], Llama3 [14], Qwen-2.5 [39]) adopt the Continual Pre-Training (CPT) strategy of concentrating high-quality data towards the end of the pre-training process. This phase is typically accompanied by learning rate decay, enabling the model to fully absorb this high-quality data. Due to significant changes in data distribution and the learning rate schedule, this approach often yields substantial improvements in metrics. Predicting a large model’s final capability based solely on its performance during a “stable” phase with consistent data distribution does not reflect its ultimate capability. Therefore, we supplement this by providing metric predictions for the high-quality CPT phase.
The relationship between model parameter scale and the volume of CPT tokens is listed in Table A1.
We conduct the same COD pipeline for CPT models. We control the data distribution of the stable and decay phases for various smaller models, as well as their token-to-parameter ratio, to be consistent with the large model targeted for prediction. The last checkpoint is used for evaluation. Based on prior clustering labels, we perform fitting, extrapolation, and mapping to obtain the predicted performance for the large model.
Table A3: Predicted vs. actual metrics for an LLM with 70B parameters after high-quality continued pretraining.
Evaluation Set Predicted Metric Actual Metric Prediction Error
MMLU-pro 61.58 59.34 2.24
BBH 84.66 85.32 0.66
TriviaQA 79.64 84.05 4.41
MBPP 73.00 73.20 0.20
AGIEval 61.66 64.18 2.52
DROP 81.65 81.39 0.26
gsm8k 86.88 91.81 4.93
MATH 49.29 52.68 3.39
Results listed in Table A3 show that the proposed COD method achieves an average prediction error of 2.33%.
We observe that MATH, GSM8k, and TriviaQA exhibit relatively large prediction errors. We hypothesize that there are two main categories of reasons for this inaccuracy:
  1. 1.
    The CPT data and the evaluation sets possess a significant correlation. For example, in math-related evaluation sets, a modest amount of training can yield substantial improvements in performance metrics. In such scenarios, the metrics for smaller models tend to show greater volatility and have inaccurate evaluations.
  2. 2.
    The CPT data exhibits inherent distribution bias, such that certain evaluation sets, such as TriviaQA, do not derive performance gains from it. This leads to potential significant fluctuations in the metrics after the CPT phase, thereby diminishing the accuracy of extrapolating to larger models.
Appendix E Additional Ablation Studies
E.1 Comparison of Clustering Methods on Extra Evaluation Sets.
We provide additional clustering evaluation results across the remaining evaluation sets in Table A4 and Table A5, which maintain consistency with the conclusions presented in the main text.
Table A4: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
Method TriviaQA AGIEval DROP MBPP
[…]
K-Means 0.4388 - 0.4572 - 0.5554 - 0.3383 -
DBScan 0.7039 6.38 0.5591 3.67 0.6651 11.08 0.5060 12.80
MeanShift 0.2521 6.77 0.2886 2.99 0.2507 11.81 0.2167 15.60
Improved-KMeans 0.1239 11.97 0.1536 7.60 0.1428 21.42 0.1667 19.40
Improved-MeanShift 0.1871 11.54 0.2100 11.50 0.1974 19.88 0.1745 21.60
Table A5: Prediction errors on advanced task benchmarks.
Method Mean Max TriviaQA AGIEval DROP MBPP
EE FE EE FE EE FE EE FE EE FE EE FE
K-Means 2.44 3.36 2.97 8.99 2.97 2.46 2.61 2.68 1.66 1.64 2.53 2.67
DBScan 3.04 2.79 6.43 4.36 1.11 0.81 6.43 6.27 3.03 2.66 1.57 1.41
MeanShift 3.21 2.34 4.18 4.90 3.64 4.90 2.63 3.23 4.18 4.00 2.40 1.22
Improved-KMeans 3.38 3.80 5.96 5.56 1.18 1.12 5.96 5.56 3.99 5.24 2.39 3.25
Improved-MeanShift 1.13 1.61 1.58 2.38 1.58 1.64 1.11 2.38 0.26 0.23 1.56 2.19
E.2 Interpolation Method
To evaluate different interpolation methods for prediction accuracy, we compared various mathematical approaches. Our baseline method uses quartic polynomial interpolation, which we compare against several alternative approaches, including Cubic spline interpolation, Cubic polynomial interpolation, and Quintic polynomial interpolation
The comparative results across different benchmarks are shown in Table A6. We report the prediction error between the real performance of a large model and the mapping result.
Table A6: Comparison of different interpolation methods across benchmarks.
Prediction Error BBH Math MMLU-pro
Cubic Spline 0.68 1.31 1.37
Cubic Polynomial 3.38 1.12 1.35
Quintic Polynomial 0.18 1.42 1.36
Quartic Polynomial 1.77 1.28 1.35
Furthermore, in Fig. A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
The prediction performance shows certain robustness across different interpolation methods. We aim to use the simplest possible interpolation function while maintaining low prediction errors. Based on the above results, we observe that the Cubic Polynomial shows larger prediction errors due to underfitting on BBH, while the Cubic Spline exhibits some overfitting. Both Quartic Polynomial and Quintic Polynomial perform well, therefore, we chose the Quartic Polynomial method as it requires fewer fitting parameters.
Figure A1: Performance mapping with different interpolation methods on the BBH evaluation set. The cubic spline is overfitted, and the cubic polynomial method is underfitted. Quartic polynomials and quintic polynomials are comparable, while a quartic polynomial has fewer parameters.
Appendix F Difficulty Distribution of Predictable Subset
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig. A2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subsets, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the MATH evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH exhibits consistent proportions of its predictable subset across varying difficulty levels, as some questions display oscillatory patterns with limited improvement, even with increased computational resources.
The proportion of the predictable subset can serve as a metric for assessing evaluation set quality. Evaluation sets with larger predictable subsets yield more reliable experimental conclusions from smaller models. When constructing evaluation sets, we recommend screening or supplementing unpredictable clusters and ensuring a minimum number of questions for each difficulty feature to reduce metric volatility.
Figure A2: Difficulty distribution comparison on a 12B model between predictable subset and full evaluation set.
Appendix G Limitations
Influence of model structure. Mixture-of-Experts (MoE) models excel in training and inference cost, and are widely used in production. In this work, we reveal the performance scaling on dense transformers, while prediction on MoE models is still underexplored. However, we believe that the proposed method is not significantly affected by the model architecture. We expect similar results for MoE models.
Category of evaluation sets. The proposed Clustering-on-Difficulty method requires a sufficient number of test cases, as too few samples can lead to unstable cluster metrics and ineffective estimation. From an evaluation set design perspective, an evaluation set with good predictive properties enables more effective generalization from small-scale to large-scale models, thus providing better guidance for model iteration.
Furthermore, for multiple-choice tasks, the model only needs to assign a higher probability to the correct option compared to others, creating a discrepancy between the answer loss and the model’s true passrate. Given that more evaluation sets are adopting Chain-of-Thoughts prompts, we have not included multiple-choice tasks that only require option selection.
Chain-of-thought performance prediction. theorem A1 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [32, 4] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.